OpenStudy (anonymous):

Simplify: ((4y^2 - 13y + 3) / (2y^2 - 5y - 12)) * ((2y^2 + 9y + 9) / (16y^2 - 1)) * ((y^2 + 3y - 28) / (y^2 - 9))

Nnesha (nnesha):

waooo

OpenStudy (anonymous):

thats for a picture of it

Nnesha (nnesha):

factor everything !!

OpenStudy (anonymous):

it is to be simplified, not factored.

Nnesha (nnesha):

\[\large\rm \frac{ 4y^2-13y+3 }{ 2y^2-5y-12 } \times \frac{ 2y^2+9y+9 }{ 16y^2-1 } \times \frac{ y^2+3y-28 }{ y^2-9 }\]

Nnesha (nnesha):

yes but you have to factor all quadratic equations

OpenStudy (anonymous):

it has been awhile since i have last did some math, or even used math in my daily life. I sadly also have a horrible memory, so i cant remember how to do this. thankfully though, i am a fast learner.

Nnesha (nnesha):

alight 4y^2 -13y +3 there are 3+ ways to factor quadratic equations easy one is to find two number if you multiply them you should get product of AC and if you add or subtract them you should get middle term

Nnesha (nnesha):

abc values are \[\large\rm Ax^2+Bx+C=0\]

OpenStudy (anonymous):

hmmmm

OpenStudy (anonymous):

maybe it is just me, but. . . i dont see how this can apply to the current equation, because the equation i posted has 3 fractions multiplying with eachother, and from what you posted. . . it is a single line with a equal sign.

Nnesha (nnesha):

okay let's do this one together \[\large\rm 4y^2 -13y +3 \]\[\large\rm Ax^2+Bx+C \] A=4 C =3 multiply AC= 12 now find two number if you multiply them you should get 12 and if you add or subtract them then you should get -13

Nnesha (nnesha):

i know we have to factor all quadratic equation one by one 4y^2 -13y +3 ^^so that equation is at the numerator of first fraction

OpenStudy (anonymous):

how does A=4 and C=3?

Nnesha (nnesha):

Quadratic equation Ax^2+Bx+C A is the leading coefficient and C is the constant term b = middle term

Nnesha (nnesha):

4y^2 -13y +3 so what is leading coefficient here and what is constant term ?

OpenStudy (anonymous):

A= 4 B=-13 c= 3?

Nnesha (nnesha):

yes that's right

Nnesha (nnesha):

multiply AC= 4 times 3 = 12 B=-13 now find two number if you multiply them you should get 12 and if you add or subtract them then you should get middle term which is -13

OpenStudy (anonymous):

sorry, my computer died. lol

OpenStudy (anonymous):

what do you mean by "find two number"?

Nnesha (nnesha):

what method did you learn 'n school ?

OpenStudy (anonymous):

i cant remember, it has been some time now... like 3-10 weeks

Nnesha (nnesha):

what two number you should multiply to get =12 1st number TIMES 2nd number = 12 what are the numbers ?

OpenStudy (anonymous):

3 & 4

Nnesha (nnesha):

yep 3 times 4 = 12 but 3+4 = -13 ?? 2 times 6 = 12 but 2 +6 isn't equal to -13 -12 times -1 = 12 correct -12 -1 = -13 ? right two numbers are -12 aand -1 when you multiply them you will get 12 and when you add them u will get -13 correct thats what we need

Nnesha (nnesha):

 carry down first and last term write those two number in middle now apply group method